OpenStudy (anonymous):

Calc III: At what point does the curve (2t^2, 1-t, 3+t^2) intersect the plane 3x - 14y + z - 10 = 0

OpenStudy (phi):

Substitute 2t^2 for x, 1-t for y and 3+t^2 for z in the equation for the plane. solve for t

OpenStudy (anonymous):

oh, easy

OpenStudy (anonymous):

thanks!

OpenStudy (phi):

I get two different points, t= 1 and t= -3

OpenStudy (anonymous):

that's the answer the book gives. out of curiosity, how did you solve the quadratic?

OpenStudy (phi):

3x - 14y + z - 10 = 0 2t^2 for x, 1-t for y and 3+t^2 for z we get 6t^2 -14(1-t) + 3+t^2 -10 = 0 7t^2 +14t -21 =0 or, dividing by 7, t^2 +2t -3=0 Now factor into (t+3)(t-1)=0 so one or the other must be zero: t+3= 0, t=-3 or t-1=0, t=1 The actual x,y,z points using 2t^2 for x, 1-t for y and 3+t^2 for z, are for t=1, (2, 0, 4) t=-3 (18, 4, 12)